# Diffusion Models as Dataset Distillation Priors

### v1

@@ 1 Introduction
Do the priors in vanilla DMs satisfy the requirements for DD? To answer this, we align the desired properties of distilled datasets with the priors captured by DMs via the original score function. From the perspective of log-likelihood estimation and evaluation metrics (e.g., FID, IS), we observe that the inherent diversity and generalization priors in vanilla DMs can yield higher-quality synthetic data. Naturally, the main challenge shifts to enhancing the representativeness of synthetic data, which is still not embodied in vanilla sampling pipelines. Previous approaches attempt to address this by imposing external representativeness constraints (Chan-Santiago et al., 2025; Chen et al., 2025). However, we argue that such constraints are unnecessary and introduce additional complexity. Thus, we raise the next question.
@@ 3.2.1 Inherent Diversity and Generalization Priors
A key question in evaluating generative models is whether they capture the full variability of the dataset (Alaa et al., 2022). We believe that DMs inherently encode diversity and generalization priors through estimating $\nabla_{\bm{x}}\log p(\bm{x})$, which aligns with the requirements of DD. This formulation compels the model to capture global manifold geometry rather than memorizing individual samples, thereby avoiding mode collapse (Thanh-Tung & Tran, 2020). Moreover, the stochastic perturbations in the forward process act as implicit regularizers, enforcing Lipschitz continuity and improving robustness to distributional shifts (Chen et al., 2024a; b). As a result, diffusion models provide a principled foundation for DD, since effective DD requires distilled data that both cover diverse modes (diversity) and faithfully approximate the original dataset distribution (generalization).
@@
We quantify these properties with likelihood-based evaluations. The negative log-likelihood (NLL) is defined as $\mathcal{L}_{\mathrm{NLL}}=-\mathbb{E}_{\bm{x}\sim p_{\rm{data}}}[\log p_{\theta}(\bm{x})]$. Identical and low NLL values on training and testing sets indicate that $p_{\theta}(\bm{x})$ converges to $p_{\rm{data}}$ instead of overfitting (see table 1). In addition, we denote entropy as $H$ and inception classifier as $\varphi$. High Inception Score (IS) indicates uniform class coverage (high $H(p_{\varphi}(y))$) and discriminative sample quality (low $H(p_{\varphi}(y|\bm{x}))$). While low Fréchet Inception Distance (FID) certifies alignment between generated and real distributions ($p_{syn}\simeq p_{\rm{data}}$). Empirical results (Dhariwal & Nichol, 2021) demonstrate that the structure-induced priors within DMs produce sufficient diversity and generalization.
@@ 3.3 Diffusion As Representativeness Prior
Let kernel function $\mathcal{K}(x,y):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be the similarity measurement which characterizes the similarity between a synthetic sample ${x}^{syn}$ and a single training sample ${x}^{train}$. We argue that the larger the similarity between the synthetic samples and the entire training set $\mathbb{E}_{\bm{x}^{train}}[\mathcal{K}(\bm{x}^{syn},\bm{x}^{train})]$, the better representativeness of $\bm{x}^{syn}$ to the raw dataset. Suppose that $\mathcal{D}_{\mathcal{K}}(x,y)$ is a distance measure induced by the kernel function $\mathcal{K}$. Typically, we expect $\mathcal{D}_{\mathcal{K}}$ to satisfy the fundamental properties of the distance measures. The following theorem demonstrates that, as long as the kernel function $\mathcal{K}$ is positive semi-definite (PSD), the induced distance $\mathcal{D}_{\mathcal{K}}$ is a well-defined distance measure.
@@
Mercer kernel allows us to quantify representativeness in RKHS, and the associated kernel-induced measure ensures the underlying optimization problem remains convex and tractable. Hence, the task reduces to identifying a suitable feature extractor $\phi$ that maps inputs into feature space, where the distance metric $d\left(\phi\left(x\right),\phi\left(y\right)\right)\propto\frac{1}{\mathcal{K}(x,y)}$. We posit that the diffusion model itself is a good feature extractor, supported by two observations: its strong image-text alignment reflects a comprehensive understanding of visual content (Yang & Wang, 2023), and its performance as a discriminative classifier exhibits high accuracy, robustness, and certified robustness (Chen et al., 2024a; b).
@@
which is referred to as energy-based guidance (Dhariwal & Nichol, 2021). Practically, we use the classifier guidance method, which employs the pre-trained diffusion itself as a training-free time-dependent classifier $\phi(x_{t})$ such that $\phi(x_{t},t)\approx\phi(x_{0})$ (Shen et al., 2024). Therefore, the reverse diffusion process with guidance is defined as:
@@
Empirically, we compare the salient features across samples using the linear kernel (Mercer kernel $\mathcal{K}(x,y)=x^{\top}y$) due to its tractability. As indicated by eq. 6, the representativeness of $\bm{x}^{syn}$ increases as the energy $\mathcal{D}_{\mathcal{K}}$ decreases. Figure 2 visualizes the representativeness of class-wise samples under different setups. The distillation performance improves on synthetic samples with higher representativeness, as reflected by the area of the sector. It is worth noting that according to eq. 4, the gradient field of diversity and generalization ($\nabla_{\bm{x}}\log p(\bm{x})$) is determined and fixed by pre-trained DMs. Therefore, the gradient field of representativeness cannot be increased indefinitely, otherwise the other priors will lose their effectiveness (see section 4.3 and section A.5.1). Hereto, we successfully distilled the prior knowledge within DMs into the synthetic dataset. We implement the guided sampling process using VP-SDE and summarize the procedure in algorithm 1. The extensive experimental results in section 4 demonstrate the validity of our “diffusion as priors” method.
@@ 4 Experiments
We evaluate DAP on ImageNet-1K and its widely used subsets (ImageNette, ImageWoof, and ImageIDC), comparing against advanced DD methods, including Minimax, D4M, IGD, and MGD3, etc. We employ two diffusion architectures, U-Net-based Stable Diffusion (SD) and Transformer-based DiT, for distillation. We also use them as baselines to demonstrate the advantage of the diffusion priors. All results are reported under either hard-label or soft-label evaluation protocols, as specified by the benchmarks. Further experimental details are provided in the section A.3.
@@ 4.1.1 Results on DiT
Dataset IPC SRe2L G-VBSM RDED Minimax DiT IGD MGD3 DAP
ImageNet-1K 10 $21.3_{\pm 0.6}$ $31.4_{\pm 0.5}$ $2.0_{\pm 0.1} 44$$4.3_{\pm 0.5}$ $39.6_{\pm 0.4}$ $45.5_{\pm 0.5}$ $\underline{45.6}_{\pm 0.8}$ $\textbf{49.1}_{\pm 1.2}$
50 $46.8_{\pm 0.2}$ $51.8_{\pm 0.4}$ $56.5_{\pm 0.1}$ $58.6_{\pm 0.3}$ $52.9_{\pm 0.6}$ $59.8_{\pm 0.3}$ $\underline{60.2}_{\pm 0.1}$ $\textbf{62.7}_{\pm 1.5}$
@@ 4.2 Analysis of Diffusion Priors
Across all benchmarks and analyses, DAP achieves competitive performance and surpasses existing DD methods. The improvements arise from the combined effect of diffusion priors: diversity and generalization priors contribute to broad coverage and cross-architecture transfer. Meanwhile, the representativeness prior enforces information alignment with the real dataset. Moreover, DAP introduces no extra training cost, which makes the approach both efficient and scalable in scenarios where deployment architectures are agnostic. We also discuss the sampling costs in section A.4.6
@@ 5 Conclusion
This paper introduces Diffusion as Priors, a framework for dataset distillation that leverages the inherent priors of diffusion models. We identify diversity, generalization, and representativeness priors in diffusion models, and demonstrate how they can be integrated to guide the generation process. Representativeness prior is further formulated through kernel–based energy guidance, enabling the sampling process to align more information with real data. Extensive experiments on ImageNet-1K and its subsets demonstrated that DAP achieves state-of-the-art results, preserves generalization under scale reduction, transfers effectively across architectures, and remains robust under domain shifts, making the approach both efficient and scalable. Future work may fall in extending diffusion priors to other powerful models (e.g., FLUX) and exploring applications beyond vision, including language, video, and multimodal datasets.
@@ A.1.1 Dataset Distillation
Systematic analysis of research in dataset distillation reveals two paradigms: a) traditional matching-based approaches focused on pixel-level optimization, and b) modern generative frameworks emphasizing distribution learning (Yu et al., 2023; Lei & Tao, 2023; Liu & Du, 2025). Traditional methods adopt an “imitation” philosophy, involving continuous pixel optimization to align model behavior, such as gradients, feature distributions, or checkpoints between synthetic and original data (Zhao & Bilen, 2021; Wang et al., 2022; Zhao et al., 2023; Deng et al., 2024). In contrast, generative frameworks prioritize improving dataset quality through fidelity and diversity metrics. These approaches extract key informational patterns from source data, enhancing the realism and generalization of distilled datasets. We will examine the related work in the following subsection.
@@ A.1.2 Generative Dataset Distillation
Diffusion-based methods leverage diffusion models to improve dataset distillation. For example, Minimax diffusion (Gu et al., 2024) fine-tunes a diffusion model with minimax criteria to boost representativeness and diversity. D4(Su et al., 2024) disentangles feature extraction and generation via Training-Time Matching (TTM) with category prototypes. IGD (Chen et al., 2025) guides the sampling process of pre-trained diffusion models using a function combining trajectory influence and diversity constraints, generating synthetic data without retraining. Additionally, MGD3 (Chan-Santiago et al., 2025) enhances diversity by identifying latent space modes and directing data toward them during sampling.
@@ A.2.2 Distance factorization
A.3 Experimental Setup
A.3.1 Datasets and benchmarks
A.3.2 Models and evaluation protocols
A.3.3 Other details
All experiments were implemented in PyTorch and conducted in a distributed setup using $8\times$ NVIDIA A40 GPUs. For fair comparison, we reproduce baseline methods under the same setup. The reported results follow these conventions: a) For DAP and reproduced baselines, we report the $\text{mean}_{\pm\text{standard deviation}}$ over three runs. b) For other methods, we report results from the original papers. c) In tables, the best result is highlighted in bold, while the second best is underlined.
A.4 Discussions
A.4.1 Compatibility of DAP
@@
IGD Influence Function
@@
Since the DAP pipeline works orthogonally with the existing approaches, it exhibits strong compatibility, allowing it to complement them without interference. Before empirical results, we analyze the sources of representativeness priors employed by different methods in table 6: Minimax introduces representativeness via fine-tuning under the supervision of the proposed training loss, D4M and MGD3 capture representativeness through clustering algorithms, IGD uses influence functions, while DAP exploits the representativeness priors embedded in diffusion models. To validate the compatibility of DAP, we incorporate it into Minimax for instance (see table 7). The addition of DAP consistently enhances the performance of distilled samples. These results indicate that DAP can serve as a versatile and modular enhancement, improving the performance of DD approaches while preserving its intrinsic advantages.
@@
A.4.2 Guidance on noisy latent
@@ A.4.3 Kernel selection
A.4.4 Cross-dataset generalization

### v2

@@ 1 Introduction
Do the priors in vanilla DMs satisfy the requirements for DD? To answer this, we align the desired properties of distilled datasets with the priors captured by DMs via the original score function. From the perspective of log-likelihood estimation and evaluation metrics (e.g., FID, IS), we observe that the inherent diversity and generalization priors in vanilla DMs can yield higher-quality synthetic data. Naturally, the main challenge shifts to enhancing the representativeness of synthetic data, which is still not embodied in vanilla sampling pipelines. Previous approaches attempt to address this by imposing external representativeness constraints (Santiago et al., 2025; Chen et al., 2025). However, we argue that such constraints are unnecessary and introduce additional complexity. Thus, we raise the next question.
@@ 2.2 Diffusion Models
2.3 Inherent Priors in Diffusion Models
A key question in evaluating generative models is whether they capture the full variability of the dataset (Alaa et al., 2022). DMs inherently encode diversity and generalization priors through estimating $\nabla_{\bm{x}}\log p(\bm{x})$, which compels the model to capture global manifold geometry rather than memorizing individual samples, thereby avoiding mode collapse (Thanh-Tung and Tran, 2020). Moreover, the stochastic perturbations in the forward process act as implicit regularizers, enforcing Lipschitz continuity and improving robustness to distributional shifts (Chen et al., 2024a; b).
In addition, let $H$ denote entropy, and $\varphi$ is an inception classifier. High Inception Score (IS) indicates uniform class coverage (high $H(p_{\varphi}(y))$) and discriminative sample quality (low $H(p_{\varphi}(y|\bm{x}))$). While low Fréchet Inception Distance (FID) certifies alignment between generated and real distributions ($p_{syn}\simeq p_{\rm{data}}$). Empirical results (Dhariwal and Nichol, 2021) demonstrate that the structure-induced priors within DMs produce sufficient diversity and generalization.
@@ 3.2.1 Inherent Diversity and Generalization Priors
As mentioned in section 2.3, diffusion models provide a principled foundation for DD, since effective DD requires distilled data that both cover diverse modes (diversity) and faithfully approximate the original dataset distribution (generalization). We quantify these properties with likelihood-based evaluations. The negative log-likelihood (NLL) is defined as $\mathcal{L}_{\mathrm{NLL}}=-\mathbb{E}_{\bm{x}\sim p_{\rm{data}}}[\log p_{\theta}(\bm{x})]$. Identical and low NLL values on training and testing sets indicate that $p_{\theta}(\bm{x})$ converges to $p_{\rm{data}}$ instead of overfitting (see table 1).
@@ 3.3 Diffusion As Representativeness Prior
To capture the representativeness prior hidden in the DMs backbone network, we require a similarity measure that quantifies how closely a synthetic sample reflects the characteristics of the real sample. Kernel function is a simple yet effective tool for defining similarity, allowing us to a) express representativeness through an induced distance and b) inject this representativeness as a differentiable energy term into the sampling process. Formally, let kernel function $\mathcal{K}(x,y):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be the smooth and differentiable similarity measurement which characterizes the similarity between a synthetic sample ${x}^{syn}$ and a single training sample ${x}^{train}$. We argue that the larger the similarity between the synthetic samples and the entire training set $\mathbb{E}_{\bm{x}^{train}}[\mathcal{K}(\bm{x}^{syn},\bm{x}^{train})]$, the better representativeness of $\bm{x}^{syn}$ to the raw dataset. Suppose that $\mathcal{D}_{\mathcal{K}}(x,y)$ is a distance measure induced by the kernel function $\mathcal{K}$. Typically, we expect $\mathcal{D}_{\mathcal{K}}$ to satisfy the fundamental properties of the distance measures. The following theorem demonstrates that, as long as the kernel function $\mathcal{K}$ is positive semi-definite (PSD), the induced distance $\mathcal{D}_{\mathcal{K}}$ is a well-defined distance measure.
@@
Mercer kernel allows us to quantify representativeness in RKHS, and the associated kernel-induced measure ensures the underlying optimization problem remains convex and tractable. Hence, the task reduces to identifying a suitable feature extractor $\phi$ that maps inputs into feature space, where the distance metric $d\left(\phi\left(x\right),\phi\left(y\right)\right)\propto\frac{1}{\mathcal{K}(x,y)}$. We posit that the diffusion model itself is a good feature extractor, supported by two observations: its strong image-text alignment reflects a comprehensive understanding of visual content (Yang and Wang, 2023), and its performance as a discriminative classifier exhibits high accuracy, robustness, and certified robustness (Chen et al., 2024a; b).
@@
which is referred to as energy-based guidance (Dhariwal and Nichol, 2021). Practically, we use the classifier guidance method, which employs the pre-trained diffusion itself as a training-free time-dependent classifier $\phi(x_{t})$ such that $\phi(x_{t},t)\approx\phi(x_{0})$ (Shen et al., 2024). Therefore, the reverse diffusion process with guidance is defined as:
@@
Empirically, we compare the salient features across samples using the linear kernel (Mercer kernel $\mathcal{K}(x,y)=x^{\top}y$) due to its tractability. As indicated by eq. 6, the representativeness of $\bm{x}^{syn}$ increases as the energy $\mathcal{D}_{\mathcal{K}}$ decreases. Figure 2 visualizes the representativeness of class-wise samples under different setups. The distillation performance improves on synthetic samples with higher representativeness, as reflected by the area of the sector. It is worth noting that according to eq. 4, the gradient field of diversity and generalization ($\nabla_{\bm{x}}\log p(\bm{x})$) is determined and fixed by pre-trained DMs. Therefore, the gradient field of representativeness cannot be increased indefinitely, otherwise the other priors will lose their effectiveness (see section 4.4 and section A.4.1).
Hereto, we successfully distilled the priors within DMs into the synthetic dataset. Specifically, Diversity prior arises from the stochasticity of diffusion trajectories where different noise initializations lead to distinct denoising paths. Generalization prior stems from the original score function $\nabla_{\bm{x}}\log p(\bm{x})$ estimated by vanilla DM. Representativeness prior is directly implemented by the guidance term in eq. 8, which guides each denosing trajectory toward gradient regions that are well represented by real data. We implement the guided sampling process using VP-SDE and summarize the procedure in algorithm 1. The extensive experimental results in section 4 demonstrate the validity of our “Diffusion As Priors (DAP)” method.
@@ 4 Experiments
We evaluate DAP on ImageNet-1K and its widely used subsets (ImageNette, ImageWoof, and ImageIDC), comparing against advanced DD methods, including Minimax, D4M, IGD, MGD3, D3HR and VLCP. We employ two diffusion architectures, U-Net-based Stable Diffusion (SD) and Transformer-based DiT, for distillation. We also use them as baselines to demonstrate the advantage of the diffusion priors. All results are reported under either hard-label or soft-label evaluation protocols, as specified by the benchmarks. Further experimental details are provided in the section 4.1.
4.1 Experimental Setup
4.1.1 Datasets and benchmarks
4.1.2 Models and evaluation protocols
4.1.3 Other details
All experiments were implemented in PyTorch and conducted with NVIDIA A40 GPUs. For fair comparison, we reproduce baseline methods under the same setup. The reported results follow these conventions: a) For DAP and reproduced baselines, we report the $\text{mean}_{\pm\text{standard deviation}}$ over three runs. b) For other methods, we report results from the original papers. c) In tables, the best result is highlighted in bold, while the second best is underlined. Practically, DAP does not require selecting any specific $x^{train|c}$ samples before sampling.
@@ 4.2.1 Results on DiT
Dataset IPC SRe2L G-VBSM RDED Minimax DiT IGD MGD3 D3HR VLCP DAP
ImageNet-1K 10 $21.3_{\pm 0.6}$ $31.4_{\pm 0.5}$ $2.0_{\pm 0.1} 44$$4.3_{\pm 0.5}$ $39.6_{\pm 0.4}$ $45.5_{\pm 0.5}$ $5.6_{\pm 0.8} 44$$4.3_{\pm 0.3}$ $\underline{46.7}_{\pm 0.4}$ $\textbf{49.1}_{\pm 1.2}$
50 $46.8_{\pm 0.2}$ $51.8_{\pm 0.4}$ $56.5_{\pm 0.1}$ $58.6_{\pm 0.3}$ $52.9_{\pm 0.6}$ $59.8_{\pm 0.3}$ ${60.2}_{\pm 0.1}$ $59.4_{\pm 0.1}$ $\underline{60.5}_{\pm 0.2}$ $\textbf{62.7}_{\pm 1.5}$
@@ 4.3 Analysis of Diffusion Priors
Across all benchmarks and analyses, DAP achieves competitive performance and surpasses existing DD methods. The improvements arise from the combined effect of diffusion priors: diversity and generalization priors contribute to broad coverage and cross-architecture transfer. Meanwhile, the representativeness prior enforces information alignment with the real dataset. Moreover, DAP introduces no extra training cost, which makes the approach both efficient and scalable in scenarios where deployment architectures are agnostic. We also discuss the sampling costs in section A.3.7.
@@ 5 Conclusion
This paper introduces Diffusion as Priors, a framework for dataset distillation that leverages the inherent priors of diffusion models. We identify diversity, generalization, and representativeness priors in diffusion models, and demonstrate how they can be integrated to guide the generation process. Representativeness prior is further formulated through kernel–based energy guidance, enabling the sampling process to align more information with real data. Extensive experiments on ImageNet-1K and its subsets demonstrated that DAP achieves state-of-the-art results, preserves generalization under scale reduction, transfers effectively across architectures, and remains robust under domain shifts, making the approach both efficient and scalable. Future work may fall in extending diffusion priors to other powerful models (e.g., FLUX, Stable Diffusion 3.5) and exploring applications beyond vision, including language, video, and multimodal datasets.
Ethics statement
This work uses only publicly available datasets, including ImageNet-1K and its subsets (ImageNette, ImageWoof, and ImageIDC). No human subjects, private data, or sensitive information are involved.
Reproducibility statement
We have taken several measures to ensure reproducibility. All theoretical results are presented with complete proofs in section A.2. Details of datasets, backbone architectures, hyper-parameters, and evaluation protocols are provided in section 4.1, while algorithms 1 and 2 specify the guided sampling procedure. Additional visualizations and ablation results are included in section A.4 to further support empirical findings. These resources ensure that all theoretical and experimental results can be independently verified.
Acknowledgments
This work was supported by the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM101), CPSF (No. 2025M771546) and NSFC (No. 625B200186).
@@ A.1.1 Dataset Distillation
Systematic analysis of research in dataset distillation reveals two paradigms: a) traditional matching-based approaches focused on pixel-level optimization, and b) modern generative frameworks emphasizing distribution learning (Yu et al., 2023; Lei and Tao, 2023; Liu and Du, 2025). Traditional methods adopt an ”imitation” philosophy, involving continuous pixel optimization to align model behavior, such as gradients, feature distributions, or checkpoints between synthetic and original data (Zhao and Bilen, 2021; Wang et al., 2022; Zhao et al., 2023; Deng et al., 2024). In contrast, generative frameworks prioritize improving dataset quality through fidelity and diversity metrics. These approaches extract key informational patterns from source data, enhancing the realism and generalization of distilled datasets. We will examine the related work in the following subsection.
@@ A.1.2 Generative Dataset Distillation
Diffusion-based methods leverage diffusion models to improve dataset distillation. For example, Minimax diffusion (Gu et al., 2024) fine-tunes a diffusion model with minimax criteria to boost representativeness and diversity. VLCP (Zou et al., 2025) constructs text prototypes to enrich the labels with semantic information and then fine-tunes the DMs with image-text pairs. D4(Su et al., 2024) disentangles feature extraction and generation via Training-Time Matching (TTM) with category prototypes. IGD (Chen et al., 2025) guides the sampling process of pre-trained diffusion models using a function combining trajectory influence and diversity constraints, generating synthetic data without retraining. Additionally, MGD3 (Santiago et al., 2025) enhances diversity by identifying latent space modes and directing data toward them during sampling. In order to enhance the objective and conditional consistency of the distillation process, CaO2 (Wang et al., 2025) employs target-guided sample selection to optimize the latent conditionally. D3HR (Zhao et al., 2025) utilizes DDIM inversion to map the image latents to the Gaussian domain, then aligns the representative latents with the high-normality Gaussian distribution with their proposed sampling scheme.
@@ A.2.2 Distance factorization
A.3 Discussions
A.3.1 Compatibility of DAP
@@
IGD Pre-trained Classifier
CaO2 Pre-trained Classifier
D3HR Statistic Metric
@@
Since the DAP pipeline works orthogonally with the existing approaches, it exhibits strong compatibility, allowing it to complement them without interference. Before empirical results, we analyze the sources of representativeness priors employed by different methods in table 6: Minimax introduces representativeness via fine-tuning under the supervision of the proposed training loss, D4M and MGD3 capture representativeness through clustering algorithms, IGD and CaO2 uses pre-trained classifier, D3HR calculates the key statistics including mean, standard deviation and skewness, while DAP exploits the representativeness priors embedded in diffusion models. To validate the compatibility of DAP, we incorporate it into Minimax for instance (see table 7). The addition of DAP consistently enhances the performance of distilled samples. These results indicate that DAP can serve as a versatile and modular enhancement, improving the performance of DD approaches while preserving its intrinsic advantages.
@@
A.3.2 Guidance on noisy latent
@@ A.3.3 Kernel selection
A.3.4 comparison with recent methods
Table 9 presents the distillation results for ImageNet subsets on DiT. CaO2 surpasses DAP in several cases, but this advantage arises from differences in evaluation protocols. The CaO2 paper reports the best accuracy across soft-label and hard-label protocols, selecting the protocol that yields a higher performance. In contrast, D3HR, VLCP, and DAP consistently adopt the hard-label protocol in this experiment. Since the soft-label protocol typically leads to higher accuracy, the results of CaO2 likely benefit from this more permissive approach. Under a consistent hard-label evaluation setting, DAP remains competitive and performs better in most cases compared to these recent methods.
Table 9: Top-1 Accuracy on ImageNette and ImageWoof. $\dagger$: The results are evaluated with hard-label protocol except for CaO2.
[…]
Nette ConvNet-6 10 $46.0_{\pm 0.5}$ $49.8_{\pm 1.1}$ $56.2_{\pm 1.3}$ - - - $\textbf{64.8}_{\pm 0.8}$ $94.3_{\pm 0.5}$
50 $71.8_{\pm 1.2}$ $70.3_{\pm 0.8}$ $74.1_{\pm 0.6}$ - - - $\textbf{82.2}_{\pm 1.6}$
100 $79.9_{\pm 0.8}$ $78.5_{\pm 0.8}$ $78.2_{\pm 0.3}$ - - - $\textbf{85.7}_{\pm 1.3}$
ResNetAP-10 10 $54.2_{\pm 1.2}$ $60.2_{\pm 0.7}$ $62.8_{\pm 0.8}$ - - $64.8_{\pm 3.6}$ $\textbf{67.8}_{\pm 1.2}$ $94.6_{\pm 0.5}$
50 $77.3_{\pm 1.0}$ $76.7_{\pm 1.0}$ $76.9_{\pm 0.5}$ - - $81.2_{\pm 0.8}$ $\textbf{82.3}_{\pm 0.7}$
100 $81.1_{\pm 0.6}$ $80.9_{\pm 0.7}$ $80.1_{\pm 1.1}$ - - - $\textbf{86.0}_{\pm 2.1}$
[…]
50 $75.8_{\pm 1.1}$ $75.0_{\pm 1.0}$ $75.2_{\pm 0.9}$ ${84.5}_{\pm 0.6}$ - - $\textbf{82.8}_{\pm 1.1}$
100 $82.0_{\pm 0.4}$ $81.5_{\pm 0.4}$ $77.8_{\pm 0.6}$ - - - $\textbf{85.5}_{\pm 1.5}$
Woof ConvNet-6 10 $25.2_{\pm 1.1}$ $27.6_{\pm 1.2}$ $32.3_{\pm 0.8}$ - - $34.8_{\pm 2.4}$ $\textbf{37.6}_{\pm 0.9}$ $85.9_{\pm 0.4}$
50 $41.9_{\pm 1.4}$ $43.8_{\pm 1.1}$ $48.5_{\pm 1.3}$ - - ${54.5}_{\pm 0.6}$ $\textbf{55.8}_{\pm 0.4}$
100 $52.3_{\pm 1.5}$ $50.1_{\pm 0.9}$ $54.2_{\pm 1.5}$ - - $\textbf{62.7}_{\pm 1.4}$ ${62.4}_{\pm 1.2}$
ResNetAP-10 10 $31.6_{\pm 0.8}$ $29.8_{\pm 1.0}$ $39.0_{\pm 0.9}$ - ${40.7}_{\pm 1.0}$ $39.5_{\pm 1.5}$ $\textbf{41.8}_{\pm 0.7}$ $87.2_{\pm 0.6}$
50 $50.1_{\pm 1.6}$ $47.8_{\pm 1.2}$ $55.8_{\pm 1.1}$ - ${59.3}_{\pm 0.4}$ $57.3_{\pm 0.5}$ $\textbf{63.3}_{\pm 0.5}$
100 $59.2_{\pm 0.9}$ $59.8_{\pm 1.3}$ $62.5_{\pm 0.9}$ - ${64.7}_{\pm 0.3}$ $65.7_{\pm 0.5}$ $\textbf{70.8}_{\pm 1.4}$
ResNet-18 10 $30.9_{\pm 1.3}$ $30.2_{\pm 0.6}$ $40.6_{\pm 0.6}$ $45.6_{\pm 1.4}$ ${39.6}_{\pm 1.0}$ $39.9_{\pm 2.6}$ $\textbf{43.9}_{\pm 0.9}$ $89.0_{\pm 0.6}$
50 $54.0_{\pm 0.8}$ $53.9_{\pm 0.7}$ $57.4_{\pm 0.7}$ $68.9_{\pm 1.1}$ ${57.6}_{\pm 0.4}$ $58.9_{\pm 1.5}$ $\textbf{63.2}_{\pm 0.7}$
[…]
A.3.5 Cross-dataset generalization